OpenStudy (anonymous):

Which of the following equations could be the result of using the comparison method to solve the system shown? x + y = 5 2x + y = 7 -x- 5 = 7 - 2x 5 -x= 7 - 2x 5 -x= 2x- 7 @johnweldon1993

OpenStudy (johnweldon1993):

so you'll want to solve for 1 variable in each equation here solve each equation for y

OpenStudy (johnweldon1993):

y = -x + 5 y = -2x + 7 now COMPARE them...(set them equal to each other) what would that look like?

OpenStudy (anonymous):

-x+5=-2x+7

OpenStudy (johnweldon1993):

which choice does that look like??

OpenStudy (anonymous):

5 -x= 2x- 7 is what i meant

OpenStudy (johnweldon1993):

same thing just rearranged :) so which choice?

OpenStudy (anonymous):

that is the choie lol

OpenStudy (johnweldon1993):

lol oh...i was waiting for like....choice 3...or something lol...regardless....yes correct :)

OpenStudy (anonymous):

Solve the following system of equations by the substitution method. x - y = 0 x - y - 2 = 0 What is the solution set? {(1, -1)} {(-1, 1)} Ø